# Decomposing Global Feature Effects Based on Feature Interactions

### v2

@@ 5.2 Theoretical Validity
Suppose $\mathds{P}_{X,Y}$ is a probability measure where $\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j},f)=0$. The type I error probability of PINT is upper bounded by $\alpha+\epsilon$, where
$\epsilon=\max_{t>0}\biggl{[}\mathds{P}\biggl{(}\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}_{n}})>t\biggr{)}-\mathds{P}\biggl{(}\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}_{n}^{\Pi}})>t\biggr{)}\biggr{]}.$ (17)
The proof can be found in Appendix D.
Intuitively, $\mathcal{D}_{n}^{\Pi}$ is a data set with no feature effects. The extra term $\epsilon$ is zero if the algorithm $\hat{f}$ is less likely to find a spurious interaction in unpermuted data compared to a case where no effects are present. Often, this is natural: most regularization techniques would nudge a learner to detect less of the spurious effect when there are other effects to focus on. However, $\epsilon$ may be strictly positive when features are strongly dependent, and the spurious effect is easier to learn than the true one. An example of such a case is given in Appendix E.1.
@@
Suppose $\mathds{P}_{X,Y}$ is a probability measure where $\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j},f)\neq 0$. If the algorithm $\hat{f}$ is such that for every $\epsilon>0$
@@
$\mathds{P}\biggl{(}\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}_{n}})>c\biggr{)}\to 1,$ (19)
the type II error probability tends to 0. The proof can be found in Appendix D.
@@ 6 Empirical Validation and Analyses of GADGET
In all our experiments and real-world applications, we control the number of partitioning steps performed by GADGET based on (1) hyperparameters typically used by decision trees, the tree depth and the minimum number of observations per leaf node, and (2) an early stop mechanism where a further split is only performed if the relative improvement of the split to be performed is at least $\gamma\in[0,1]$ times the total relative interaction-related heterogeneity reduction of the previous split: $\gamma\times\frac{\sum_{j\in S}\left(\mathcal{R}(\mathcal{A}_{P},\tilde{\mathbf{x}}_{j})\right)-\mathcal{I}(\hat{t},\hat{z})}{\sum_{j\in S}\left(\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})\right)}$. The higher we choose $\gamma$, the fewer partitioning steps will be performed. See Appendix C for more details and recommendations.
@@ Results.
Figure 4 shows that independent of the model or correlation degree, $\mathbf{x}_{2}$ has (correctly) never been considered as split feature. For correlation strengths $\rho_{13}$ between 00 and $0.7$, $\mathbf{x}_{3}$ is always chosen as the only split feature, with $I_{3}$ taking values between $0.75$ and $1$ and thus reducing most of the interaction-related heterogeneity of all features with one split. Thus, for low to medium correlations, there are only minor differences between the various feature effect methods and models. However, the observed behavior changes substantially for $\rho_{13}=0.9$. For the correctly specified GAM, we still receive fairly consistent results, apart from one repetition where $\mathbf{x}_{1}$ is chosen as the split feature when PD is used. In contrast, there is more variation of $I_{3}$ for the NN, especially when SD is used. For SD, $\mathbf{x}_{1}$ is chosen once for splitting, and for PD, this is also the case for $30\%$ of all repetitions (see Table 1). Using ALE, GADGET always correctly performs one split with respect to $\mathbf{x}_{3}$. Additionally, GADGET performs a second split once when PD is used and 7 times when SD is used. Thus, for high correlations between features, we already observe in this very simple setting that the extrapolation problem influences the splitting within GADGET for effect methods based on marginal distributions, while methods based on conditional distributions such as ALE are less affected. This is particularly relevant for learners that model very locally (e.g. NNs) and, thus, tend to have wiggly prediction functions and oscillate in extrapolating regions. This is also notable in the split value range, which increases for the NN in the case of high correlations for PD and SD but not for ALE (see Table 1).
@@ 6.2 Influence of Higher-Order Interaction Effects
While SD puts less weight on interactions with increasing order, all interactions (regardless of the order) receive the same weight in PD and ALE. Hence, when using GADGET-SD, we may not be able to detect high order interactions—especially when using the approximation without recalculation after each split (see Section 4.5 and Appendix B.3). However, it should be more likely to detect these interactions with PD and ALE.
@@ 6.3 Influence of Spurious Interactions
We reuse the simulation example described in Section 5: Consider four features with $X_{1},X_{2},X_{4}\sim U(-1,1)$ and $X_{3}=X_{2}+\epsilon$ with $\epsilon\sim N(0,0.09)$. For $30$ repetitions, we draw $n=\{300,500\}$ observations and create the dependent variable, including a potential spurious interaction between $\mathbf{x}_{1}$ and $\mathbf{x}_{3}$ (as $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ interact, but $\mathbf{x}_{2}$ and $\mathbf{x}_{3}$ are correlated) [^9] While the true functional relationship is here defined without noise to purely measure the effect of the spurious interaction, we analyze the effect of adding noise in Appendix E.2: $\mathbf{y}=\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{x}_{3}-2\mathbf{x}_{1}\mathbf{x}_{2}$. For each sample size $n$ and each repetition, we fit an SVM with an RBF kernel, cost parameter $C=1$, and choose the inverse kernel width based on the data. The performance of the model, measured by the mean (standard deviation) of the MSE on a test set of size 100000 across all repetitions, is $0.028(0.010)$ for $n=300$ and $0.027(0.010)$ for $n=500$. We calculate PINT using PD, ALE, and SD for each repetition and sample size with $s=100$ and $\alpha=0.05$ by approximating the null distribution as described in Section 5. We apply GADGET using PD, ALE, and SD with recalculation, where $S=Z$ is based on the feature subset chosen by PINT for each method. We compare these results by considering all features as features of interest and potential split features (i.e. $S=Z=\{1,\ldots,p\}$). We use a maximum tree depth of $6$, minimum number of observations of $40$, and $\gamma=0.15$ as stop criteria.
@@ Results.
7 Empirical Validation of the PINT test as Hypothesis Test
We now empirically validate PINT as a nonparametric test to detect interactions.
@@ 7 Empirical Validation of the PINT test as Hypothesis Test
It can be shown empirically that the PINT test is faithful w.r.t. the type I error and the power of the test as already validated theoretically in Section 5.
@@ Results.
The rejection rate of all four features and all effect methods for $\beta=0$ align closely with the significance level $\alpha=0.05$ (see Figure 8). While the power of the test approaches $1$ for increasing effect sizes of the interaction between $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ for all three methods, the type I error represented by the rejection rates of $\mathbf{x}_{3}$ and $\mathbf{x}_{4}$ goes to 00. This trend can be explained as follows: The higher $\beta$, the higher the influence of $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ and thus the more the model focuses on these two features and ignores the two non-influential features. This can also be seen on the p-value distributions, which, for non-influential features such as $\mathbf{x}_{4}$, becomes more skewed towards $1$ the higher the interaction effect, as illustrated in Figure 9. Thus, in the most extreme case, the model disregards the non-influential features entirely, resulting in their local effects’ heterogeneity being exactly 00. However, when permuting $\mathbf{y}$ and refitting the model, random effects may be learned for $\mathbf{x}_{3}$ and $\mathbf{x}_{4}$. Hence, the calculated values under the null might be higher than the test statistic value, causing the observed artifact in Figure 8 where $H_{0}$ is never falsely rejected instead of in $5\%$ of cases.
More diverse and complex settings that confirm these observations and empirically validate their correctness are discussed in Appendix E.2.
Figure 8: Rejection rates of PINT for $1000$ repetitions for all three feature effect methods and for varying interaction effect sizes $\beta$. For $\beta=0$, all rejection rates correspond to type I errors. For $\beta>0$, the rejection rates of $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ correspond to the power of the test, those of $\mathbf{x}_{3}$ and $\mathbf{x}_{4}$ to type I errors. The red dashed line represents the significance level of $\alpha=0.05$. Where not visible, the green line ($\mathbf{x}_{1}$) is below the purple line ($\mathbf{x}_{2}$).
@@ 8.1 COMPAS Data Set
Since the features do not show high correlations, we use ICE and PD for our analysis. [^11] We obtained similar results by using SD instead of PD within GADGET, see Appendix F. Figure 10 shows that the average predicted risk visibly decreases with age, while the predicted risk first increases steeply with the number of previous crimes to 10 and then decreases slightly for higher values. The PD values for the type of crime do not differ substantially. When considering the two “inadmissible” features, there is on average a slightly higher risk of recidivism for African-American versus Caucasian and female versus male defendants. For all features, we can observe highly differing local effects. Particularly, the heterogeneous ICE curves for age and number of prior crimes indicate feature interactions.
@@
The effect plots for the four resulting regions are shown in Figure 11. GADGET performed the first split according to age and the splits on the second depth according to the number of prior crimes. The total reduction in interaction-related heterogeneity is $R^{2}_{Tot}=0.86$. The highest reduction in heterogeneity is given by age and the number of prior crimes, which interact the most (see Figure 11). For defendants around 20 years of age, the predicted risk tends to be high. However, for defendants with a small number of prior crimes, the predicted risk decreases very quickly with increasing age, reaching a low risk and remaining so for people older than 32 years. In contrast, for defendants with more than four prior crimes, the predicted risk decreases only slowly with increasing age. The regional feature effects of the number of prior crimes show that the interaction-related heterogeneity is small for the subgroup of younger people, while some heterogeneity still remains for defendants older than 32 years of age. Furthermore, the interaction-related heterogeneity of the three binary features (ethnicity, gender, and severity of crime) was reduced, indicating an interaction between each of them and age as well as the number of previous crimes. Although the effect on the predicted risk only differs slightly between the categories of the three binary features for older defendants with a lower number of prior crimes and for younger defendants with a high number of prior crimes, greater differences were observed for the other two subgroups. The overall difference in predicted risk for the two inadmissible features of ethnicity and gender seems to be especially high for people over 32 years with a higher number of prior crimes, as well as for people younger than 32 with a lower number of prior crimes. These differences in the predicted risk between subgroups of defendants might be due to a potentially learned bias regarding the ethnicity and gender of the defendant and thus potentially resulting in a more severe predicted risk of recidivism for some defendants than for others.
@@
Figure 11: Regional PD plots for all features of the COMPAS application after applying GADGET. The grey areas of the numeric features and the error bars of the categorical features indicate the $95\%$ interaction-related heterogeneity interval as defined in Appendix B.4.
@@ 8.2 Bikesharing Data Set
We use the bikesharing data set (James et al., 2022), to predict the hourly counts of rented bikes from the Capital bikeshare system for the years 2011 and 2012. We fit a RF with 10 seasonal and weather features: the day and hour the bike was rented, if the current day is a working day, the season, number of casual bikers, normalized temperature, perceived temperature, wind speed, humidity, and weather situation (categorical, e.g., “clear”).
@@
Figure 12: Global (left) and regional ALE plots after applying GADGET for feature hour using $S=Z=\{hr,workingday\}$ (middle) and using $S=Z=\{hr,workingday,temp\}$ (right) of the bikesharing data. The upper plots show the interaction-related heterogeneity as defined in Appendix B.4.
@@ 9 High-Dimensional Real-World Application with Filtering Procedure
To decide on the cutoff point of the risk values to group features into one of the two groups: heterogeneous or homogeneous local effects, we visualize the risk values calculated by Eq. (7) in decreasing order by an elbow graph as illustrated in Figure 13. The cutoff point is chosen rather conservatively as the last substantial drop in risk before all remaining values are close to 00.131313This cut-off point can also be chosen in a data-driven way, e.g., by the percentage of explained variance similarly to principal component analysis. Figure 14 visualizes the respective mean-centered ICE curves for the two features closest to the cutoff point and compares them to the mean-centered ICE curves of the features with the highest and smallest risk values according to Figure 13. This first filtering step leaves $30$ features in the group of heterogeneous local effects.
@@
Figure 14: Mean-centered ICE curves for the feature with the highest (left) and lowest (right) risk values and the two features closest to the cut-off point (middle).
We then apply PINT as a second filtering step, with PD, $s=200$ permutations, and $\alpha=0.05$ as described in Section 5. $12$ out of the $30$ features are detected as significant. We apply GADGET-PD with a relative improvement of $\gamma=0.25$, a maximum tree depth of $5$ and minimum number of observations per node of $100$ as early stopping criteria. [^14] See Appendix C for general recommendations on specifying these hyperparameters. We define $S=Z$ as the $12$ obtained features. We obtain $5$ final regions when GADGET-PD is applied. With $R^{2}_{Tot}=0.67$, we removed $67\%$ of the interaction-related heterogeneity within the final regions of all $12$ features in $S$. The visualizations and interpretations of the regional effect plots are provided in Appendix G.

### v3

@@ 5.2 Theoretical Properties
Suppose $\mathds{P}_{X,Y}$ is a probability measure where $\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j},f)=0$. The type I error probability of PINT is upper bounded by $\alpha+\epsilon$, where
$\epsilon=\max_{t>0}\biggl{[}\mathds{P}\biggl{(}\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}_{n}})>t\biggr{)}-\mathds{P}\biggl{(}\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}_{n}^{\Pi}})>t\biggr{)}\biggr{]}.$ (17)
The proof can be found in Appendix E.
Intuitively, $\mathcal{D}_{n}^{\Pi}$ is a data set with no feature effects. The extra term $\epsilon$ is zero if the algorithm $\hat{f}$ is less likely to find a spurious interaction in unpermuted data compared to a case where no effects are present. Often, this is natural: most learners detect less of the spurious effect when there are other effects to focus on. In this case, the PINT procedure is a valid statistical test. In fact, we conjecture that the test is conservative (i.e., type I error is strictly smaller than $\alpha$) in most cases, which is in line with our empirical observations (see Section 7 and Appendix F). However, $\epsilon$ may be strictly positive when features are strongly dependent, and the spurious effect is easier to learn than the true one. An example of such a case is given in Appendix F.1.
@@
Suppose $\mathds{P}_{X,Y}$ is a probability measure where $\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j},f)\neq 0$. If the algorithm $\hat{f}$ is such that for every $\epsilon>0$
@@
$\mathds{P}\biggl{(}\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}_{n}})>c\biggr{)}\to 1,$ (19)
the type II error probability tends to 0. The proof can be found in Appendix E.
@@ 6 Empirical Validation and Analyses of GADGET
In all our experiments and real-world applications, we control the number of partitioning steps performed by GADGET based on (1) the tree depth and the minimum number of observations per leaf node which are hyperparameters typically used by decision trees, and (2) an early stopping mechanism where a further split is only performed if the relative improvement of the split to be performed is at least $\gamma\in[0,1]$ times the total relative interaction-related heterogeneity reduction of the previous split: $\gamma\times\frac{\sum_{j\in S}\left(\mathcal{R}_{j}(\mathcal{A}_{P},\tilde{\mathbf{x}}_{j})\right)-\mathcal{I}(\hat{t},\hat{z})}{\sum_{j\in S}\left(\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j})\right)}$. The higher we choose $\gamma$, the fewer partitioning steps will be performed. See Appendix D for more details and recommendations.
@@ Results.
Figure 5 shows that independent of the model or correlation degree, $\mathbf{x}_{2}$ has (correctly) never been considered as split feature. For correlation strengths $\rho_{13}$ between 00 and $0.7$, $\mathbf{x}_{3}$ is always chosen as the only split feature, with $I_{3}$ taking values between $0.75$ and $1$ and thus reducing most of the interaction-related heterogeneity of all features with one split. Thus, for low to medium correlations, there are only minor differences between the various feature effect methods and models. However, the observed behavior changes substantially for $\rho_{13}=0.9$. For the correctly specified GAM, we still receive fairly consistent results, apart from one repetition where $\mathbf{x}_{1}$ is chosen as the split feature when PD is used. In contrast, there is more variation of $I_{3}$ for the NN. For SD, $\mathbf{x}_{1}$ is chosen once for splitting, and for PD, this is also the case for $30\%$ of all repetitions (see Table 1). Using ALE, GADGET always correctly performs one split with respect to $\mathbf{x}_{3}$. Additionally, GADGET performs a second split once when PD is used and 7 times when SD is used. Thus, for high correlations between features, we already observe in this very simple setting that the extrapolation problem influences the splitting within GADGET for effect methods based on marginal distributions, while methods based on conditional distributions such as ALE are less affected. This is particularly relevant for learners that model very locally (e.g. NNs) and, thus, tend to have wiggly prediction functions and oscillate in extrapolating regions. This is also notable in the split value range, which increases for the NN in the case of high correlations for PD and SD but not for ALE (see Table 1).
@@ 6.2 Influence of Higher-Order Interaction Effects
While SD puts less weight on interactions with increasing order, all interactions (regardless of the order) receive the same weight in PD and ALE. Hence, when using GADGET-SD, we may not be able to detect high order interactions—especially when using the approximation without recalculation after each split (see Section 4.5 and Appendix C.3). However, it should be more likely to detect these interactions with PD and ALE.
@@ 6.3 Influence of Spurious Interactions
We reuse the simulation example described in Section 5: Consider four features with $X_{1},X_{2},X_{4}\sim U(-1,1)$ and $X_{3}=X_{2}+\epsilon$ with $\epsilon\sim N(0,0.09)$. For $30$ repetitions, we draw $n=\{300,500\}$ observations and create the dependent variable, including a potential spurious interaction between $\mathbf{x}_{1}$ and $\mathbf{x}_{3}$ (as $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ interact, but $\mathbf{x}_{2}$ and $\mathbf{x}_{3}$ are correlated) [^10] While the true functional relationship is here defined without noise to purely measure the effect of the spurious interaction, we analyze the effect of adding noise in Appendix F.2: $\mathbf{y}=\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{x}_{3}-2\mathbf{x}_{1}\mathbf{x}_{2}$. For each sample size $n$ and each repetition, we fit an SVM with an RBF kernel, cost parameter $C=1$, and choose the inverse kernel width based on the data. The performance of the model, measured by the mean (standard deviation) of the MSE on a test set of size 100000 across all repetitions, is $0.028\ (0.010)$ for $n=300$ and $0.027\ (0.010)$ for $n=500$. We calculate PINT using PD, ALE, and SD for each repetition and sample size with $s=100$ and $\alpha=0.05$ by approximating the null distribution as described in Section 5. We apply GADGET using PD, ALE, and SD with recalculation, where $S=Z$ is based on the feature subset chosen by PINT for each method. We compare these results by considering all features as features of interest and potential split features (i.e. $S=Z=\{1,\ldots,p\}$). We use a maximum tree depth of $6$, minimum number of observations of $40$, and $\gamma=0.15$ as stop criteria.
@@ Results.
7 Empirical Validation of the PINT Procedure
We now empirically validate PINT to detect interactions.
@@ 7 Empirical Validation of the PINT Procedure
It can be shown empirically that the PINT procedure is faithful w.r.t. the type I error and the power of the test as already validated theoretically in Section 5.
@@ Results.
The rejection rate of all four features and all effect methods for $\beta=0$ align closely with the significance level $\alpha=0.05$ (see Figure 9). While the power of the test approaches $1$ for increasing effect sizes of the interaction between $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ for all three methods, the type I error represented by the rejection rates of $\mathbf{x}_{3}$ and $\mathbf{x}_{4}$ goes to 00. This trend can be explained as follows: The higher $\beta$, the higher the influence of $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ and thus the more the model focuses on these two features and ignores the two non-influential features. This can also be seen on the p-value distributions, which, for non-influential features such as $\mathbf{x}_{4}$, becomes more skewed towards $1$ the higher the interaction effect, as illustrated in Figure 10. Thus, in the most extreme case, the model disregards the non-influential features entirely, resulting in their local effects’ heterogeneity being exactly 00. However, when permuting $\mathbf{y}$ and refitting the model, random effects may be learned for $\mathbf{x}_{3}$ and $\mathbf{x}_{4}$. Hence, the calculated values under the null distribution might be higher than the actual risk value, causing the observed artifact in Figure 9 where the type I error is 0 instead of showing a false rejection rate of $5\%$.
More diverse and complex settings that confirm these observations and empirically validate their correctness are discussed in Appendix F.2.
Figure 9: Rejection rates of PINT for $1000$ repetitions for PD, ALE, and SD and for varying interaction effect sizes $\beta$. For $\beta=0$, rejection rates correspond to type I errors. For $\beta>0$, rejection rates of $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ correspond to the power of the test, those of $\mathbf{x}_{3}$ and $\mathbf{x}_{4}$ to type I errors. The red dashed line represents the significance level of $\alpha=0.05$. Where not visible, the green line ($\mathbf{x}_{1}$) is below the purple line ($\mathbf{x}_{2}$).
@@ 8.1 COMPAS Data Set
Since the features do not show high correlations, we use ICE and PD for our analysis. [^12] We obtained similar results by using SD instead of PD within GADGET, see Appendix G. Figure 11 shows that the average predicted risk visibly decreases with age, while the predicted risk first increases steeply with the number of previous crimes up to 10 and then decreases slightly for higher values. The PD values for the type of crime do not differ substantially. When considering the two “inadmissible” features, there is on average a slightly higher risk of recidivism for African-American versus Caucasian and female versus male defendants. For all features, we can observe highly differing local effects. Particularly, the heterogeneous ICE curves for age and number of prior crimes indicate feature interactions.
@@
The effect plots for the four resulting regions are shown in Figure 12. GADGET performed the first split according to age and the splits on the second depth according to the number of prior crimes. The total reduction in interaction-related heterogeneity is $R^{2}_{Tot}=0.86$. The highest reduction in heterogeneity is given by age and the number of prior crimes, which interact the most (see Figure 12). For defendants around 20 years of age, the predicted risk tends to be high. However, for defendants with a small number of prior crimes, the predicted risk decreases very quickly with increasing age, reaching a low risk and remaining so for people older than 32 years. In contrast, for defendants with more than four prior crimes, the predicted risk decreases only slowly with increasing age. The regional feature effects of the number of prior crimes show that the interaction-related heterogeneity is small for the subgroup of younger people, while some heterogeneity still remains for defendants older than 32 years of age. Furthermore, the interaction-related heterogeneity of the three binary features (ethnicity, gender, and severity of crime) was reduced, indicating an interaction between each of them and age as well as the number of prior crimes. Although the effect on the predicted risk only differs slightly between the categories of the three binary features for older defendants with a lower number of prior crimes and for younger defendants with a high number of prior crimes, greater differences were observed for the other two subgroups. The overall difference in predicted risk for the two inadmissible features of ethnicity and gender seems to be especially high for people over 32 years with a higher number of prior crimes, as well as for people younger than 32 with a lower number of prior crimes. These differences in the predicted risk between subgroups of defendants might be due to a potentially learned bias regarding the ethnicity and gender of the defendant and thus potentially resulting in a more severe predicted risk of recidivism for some defendants than for others.
@@
Figure 12: Regional PD plots for features of the COMPAS application after applying GADGET. Grey areas of numeric features and error bars of categorical features indicate the $95\%$ interaction-related heterogeneity interval (see Appendix C.4).
@@ 8.2 Bikesharing Data Set
We use the bikesharing data set (James et al., 2022), to predict the hourly counts of rented bikes from the Capital bikeshare system for the years 2011 and 2012. We fit an RF with 10 seasonal and weather features: the day and hour the bike was rented, if the current day is a working day, the season, number of casual bikers, normalized temperature, perceived temperature, wind speed, humidity, and weather situation (categorical, e.g., “clear”).
@@
Figure 13: Global (left) and regional ALE plots after applying GADGET for feature hour using $S=Z=\{hr,workingday\}$ (middle) and using $S=Z=\{hr,workingday,temp\}$ (right) of the bikesharing data. The upper plots show the interaction-related heterogeneity as defined in Appendix C.4.
@@ 9 High-Dimensional Real-World Application with Filtering Procedure
To decide on the cutoff point of the risk values to group features into one of the two groups: heterogeneous or homogeneous local effects, we visualize the risk values calculated by Eq. (7) in decreasing order by an elbow graph as illustrated in Figure 14. The cutoff point is chosen rather conservatively as the last substantial drop in risk before all remaining values are close to 00.141414This cutoff point can also be chosen in a data-driven way, e.g., by the percentage of explained variance similarly to principal component analysis. Figure 15 visualizes the respective mean-centered ICE curves for the two features closest to the cutoff point and compares them to the mean-centered ICE curves of the features with the highest and smallest risk values according to Figure 14. This first filtering step leaves $30$ features in the group of heterogeneous local effects.
@@
Figure 15: Mean-centered ICE curves for the feature with the highest (left) and lowest (right) risk values and the two features closest to the cutoff point (middle).
We then apply PINT as a second filtering step, with PD, $s=200$ permutations, and $\alpha=0.05$ as described in Section 5. $12$ out of the $30$ features are detected as significant. We apply GADGET-PD with a relative improvement of $\gamma=0.25$, a maximum tree depth of $5$ and minimum number of observations per node of $100$ as early stopping criteria. [^15] See Appendix D for general recommendations on specifying these hyperparameters. We define $S=Z$ as the $12$ obtained features. We obtain $5$ final regions when GADGET-PD is applied. With $R^{2}_{Tot}=0.67$, we removed $67\%$ of the interaction-related heterogeneity within the final regions of all $12$ features in $S$. The visualizations and interpretations of the regional effect plots are provided in Appendix H.